FIND THE SURFACE AREA OF THE OBJECT OBTAINED BY ROTATING Y=4+3x^2, 1<=x<=2 ABOUT THE Y-AXIS


1
Expert's answer
2021-01-14T18:55:03-0500

FIND THE SURFACE AREA OF THE OBJECT OBTAINED BY ROTATING Y=4+3x2, 1<=x<=2 ABOUT THE Y-AXIS

"Surface" Integral = "2\pi \int_{1}^{2} x \sqrt{1 + [f'(x)]^2}dx"

"f(x) = y=4 + 3x^2"

"f'(x)=6x"


Surface Area = "2\pi \int_{1}^{2} x \sqrt{1 + (6x)^2}dx"


"2\pi \int_{1}^{2} x \sqrt{1 + 36x^2}dx"


let u = "1 + 36x^2" , "{du \over dx} = 72x" , "dx = {du \over 72x}"


Surface Area "=2\pi \int_{1}^{2} {x \over 72x} \sqrt{u}du"


"={2\pi \over 72} \int_{1}^{2} \sqrt{u}du"


"={\pi \over 36} .{2\sqrt{u}^3 \over 3} |_{1}^{2}"

"= {\pi\sqrt{u}^3 \over54} |_{1}^{2}"


"= {\pi(\sqrt{1 + 36x^2})^3 \over54} |_{1}^{2}"


"=[ {\pi(\sqrt{1 + 36(2)^2})^3 \over54}] - [{\pi(\sqrt{1 + 36(1)^2})^3 \over54} ]"

"=[ {\pi(\sqrt{145})^3 \over54}] - [{\pi(\sqrt{37})^3 \over54} ]"



"= {\pi \over54} [ \sqrt{(145)^3}-{\sqrt{(37)^3 }}]"

Surface Area "= {\pi \over54} [145 \sqrt{145}-37{\sqrt{37 }}]"